OpenStudy (scorcher219396):

Differentiate using the chain rule: (xe^-x)^2

OpenStudy (scorcher219396):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

well do you know how to use chain rule?

OpenStudy (scorcher219396):

I do, im just having a hard time trying to figure out what the inner and outer function are here

OpenStudy (scorcher219396):

chain rule says for f composition g, F'(x)=f'(g(x))(g'x)

OpenStudy (xapproachesinfinity):

\(\Large \tt\color{blueviolet}{f(x)=(xe^x-x)^2}\) the inner function is xex-x the outer let it be x^2

OpenStudy (xapproachesinfinity):

So \(\Large \tt\color{blueviolet}{f'(x)=2(xe^x-x)\times(xe^x-x)'}\)

OpenStudy (scorcher219396):

Im confused as to how you got the first part

OpenStudy (xapproachesinfinity):

well do it like this set f(x)=x^2 g(x)=xe^x-x okay

OpenStudy (scorcher219396):

How did you get g(x)?

OpenStudy (xapproachesinfinity):

now do f'(x)=2x correct?

OpenStudy (scorcher219396):

Since the original problem is\[y=xe ^{-x ^{2}}\]

OpenStudy (xapproachesinfinity):

it is the inner function i just set it equal to g(x) to make this clear

OpenStudy (scorcher219396):

My bad, sorry i typed the original function wrong

OpenStudy (xapproachesinfinity):

oh sorry is though is was \(\Large \tt\color{blueviolet}{xe^x-x}\)

OpenStudy (xapproachesinfinity):

alright let's go back to that function then \(\Large \tt\color{blueviolet}{F(x)=(xe^{-x}) ^2}\)

OpenStudy (xapproachesinfinity):

the inner function is \(xe^{-x}\) the outer ix \(x^2\)

OpenStudy (scorcher219396):

But there are no parentheses around the whole thing

OpenStudy (scorcher219396):

That's why im getting stuck

OpenStudy (xapproachesinfinity):

so what is like \(\Large \tt\color{blueviolet}{xe^{-2x}?}\)

OpenStudy (xapproachesinfinity):

i thought you put parenthesis in your post

OpenStudy (scorcher219396):

\[y=xe ^{-x ^{2}}\]

OpenStudy (xapproachesinfinity):

oh i see! you first need to do product rule

OpenStudy (scorcher219396):

Oh ok that makes sense

OpenStudy (xapproachesinfinity):

\(\Large \tt\color{blueviolet}{(xe^{-x^2})'=x'e^{-x^2}+x(e^{-x^2})'}\)

OpenStudy (xapproachesinfinity):

x'=1 you just need to do chain rule to the last term. yes?

OpenStudy (scorcher219396):

Right, which would be \[2e^{-x}(-e^{-x})\]? Then it would be \[-2e^{-2x}\] ?

OpenStudy (xapproachesinfinity):

hmm not really set f(x)=e^x ang g(x)=-x^2 and see

OpenStudy (xapproachesinfinity):

it would be \(\Large \tt\color{blueviolet}{-2xe^{-x^2}}\) agree?

OpenStudy (scorcher219396):

Oh ok yep that makes more sense

OpenStudy (xapproachesinfinity):

now you just need to clean that thing and factor exp(-x^2) and that's all

OpenStudy (xapproachesinfinity):

remember you sill have x attached to (e^{-x^2})

OpenStudy (xapproachesinfinity):

you get it?